Very confused on problem...I would greatly appreciate your help and time...thanks a bunch
A set of 50 data values has a mean of 28 and a variance of 4.
I. Find the standard score (z) for a data value = 31.
II. Find the probability of a data value > 31.
III. Find the probability of a data value < 31.
Show all work.

Answer by stanbon(75887)

A set of 50 data values has a mean of 28 and a variance of 4.
I. Find the standard score (z) for a data value = 31.
z(31) = (31-28)/4 = 3/4
----------------------
II. Find the probability of a data value > 31.
z(31) = 3/4
P(x > 31) = P(z > 3/4) = 0.2266
-----------------------
III. Find the probability of a data value < 31.
P(x < 31) = 1 - 0.2266 = 0.7734
==================================
Cheers,
Stan H.
===========